# Practical Principles for AI Cost and Compute Accounting

### v3

@@ Abstract
Policymakers increasingly use development cost and compute as proxies for AI capabilities and risks. Recent laws have introduced regulatory requirements that are contingent on specific thresholds. However, technical ambiguities in how to perform this accounting create loopholes that can undermine regulatory effectiveness. We propose seven principles for designing AI cost and compute accounting standards that (1) reduce opportunities for strategic gaming, (2) avoid disincentivizing responsible risk mitigation, and (3) enable consistent implementation across companies and jurisdictions.
@@ 1 Introduction
As AI systems become more capable, policymakers face a mounting challenge: identifying which AI systems warrant heightened oversight. Recent laws and governance frameworks have approached this challenge by proposing requirements contingent on development costs and computational resources (e.g., European Parliament and Council, 2024; California SB1047, 2024; Li et al., 2025; U.S. Bureau of Industry and Security, 2025). For example, the now rescinded 2025 Framework for Artificial Intelligence Diffusion, issued by the U.S. Bureau of Industry and Security, sought to control the export of AI model weights for systems whose development exceeded $10^{26}$ floating point operations (FLOP) in training compute (U.S. Bureau of Industry and Security, 2025).
Development costs and compute are compelling metrics for use in AI governance (Li et al., 2025) because they “correlate with capabilities and risks, are quantifiable, can be measured early in the AI lifecycle, and can be verified by external actors” (Heim & Koessler, 2024). Cost and compute thresholds also allow regulators to focus oversight on the most advanced AI models while avoiding unnecessary burdens on smaller developers. However, their practicality as regulatory tools depends on establishing clear accounting standards. This requires resolving several technical ambiguities about what should be counted (Hooker, 2024; Heim & Koessler, 2024; Reuel et al., 2024).

### v4

@@ Abstract
Policymakers increasingly use development cost and compute as proxies for AI capabilities and risks. Recent laws have introduced regulatory requirements for models or developers that are contingent on specific thresholds. However, technical ambiguities in how to perform this accounting create loopholes that can undermine regulatory effectiveness. We propose seven principles for designing AI cost and compute accounting standards that (1) reduce opportunities for strategic gaming, (2) avoid disincentivizing responsible risk mitigation, and (3) enable consistent implementation across companies and jurisdictions.
@@ 1 Introduction
As AI systems become more capable, policymakers face a mounting challenge: identifying which AI systems warrant heightened oversight. Recent laws, draft laws, and governance frameworks have approached this challenge by proposing requirements contingent on development costs and computational resources (e.g., European Parliament and Council, 2024; California SB1047, 2024; Li et al., 2025; U.S. Bureau of Industry and Security, 2025; New York A6453A, 2025; California SB 53, 2025; Michigan HB 4668, 2025). For example, Article 51 of the EU AI Act states “A general-purpose AI model shall be presumed to have high impact capabilities…when the cumulative amount of computation used for its training measured in floating point operations is greater than $10^{25}$.”
Development costs and compute are compelling metrics for use in AI governance (Li et al., 2025) because they “correlate with capabilities and risks, are quantifiable, can be measured early in the AI lifecycle, and can be verified by external actors” (Heim & Koessler, 2024). Cost and compute thresholds also allow regulators to focus oversight on the most advanced AI developers while avoiding unnecessary burdens on smaller developers. However, their practicality as regulatory tools depends on establishing clear accounting standards. This requires resolving several technical ambiguities about what should be counted (Hooker, 2024; Heim & Koessler, 2024; Reuel et al., 2024).